Arizona:

hellllllp (9/3/2)/9*6 show your work i know the answer is 1 by a calculator but i cant figure out how to show my work

Arizona:

@hannahj @midnight97

Arizona:

@toga please someone help this is due tomarrow

XxDevilxX:

first you want to do 9/3, then you will divide the answer of 9/3 by 2, then mulitply the 9 and the 6 in the bottom of the fraction and finally you will divide whatever the answer is to the top by whatever the answer is to the bottom. Dont know if that makes any sense or helps but I hope it does.

XxDevilxX:

I also could be wrong, im not sure. i think thats how you do it though. im like 95% sure.

toga:

First, we need to follow the order of operations, which is parentheses, multiplication and division from left to right, and then addition and subtraction from left to right. 1. Start with the parentheses: 9/3/2 = 3 2. Now we have: 3)/9*6 3. Next, we do the multiplication and division from left to right: 3) / 9 = 1/3 1/3 * 6 = 2 Therefore, the answer is 2.

XxDevilxX:

ahh seems like i might have been wrong i'd still with Toga's answer lol

XxDevilxX:

stick*

hannahj:

dont copy my name I dont feel like answering. ;-;

Arizona:

thanks for the help

toga:

@arizona wrote:
thanks for the help
np

XxDevilxX:

@arizona wrote:
thanks for the help
i didnt really help but i tried sooo like np lol